• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name V. Sulkosky

    Log entry time 10:00:34 on February 04, 2009

    Entry number 258904

    keyword=Gas Cerenkov ADC #6 Status

    I have looked at the data taken over night after I switched ADC delay
    cables for ADC #6 for the left gas Cerenkov. The amplitude has been
    stable overnight with no changes.

    Figure 1 shows the pedestal subtracted amplitude for run 5697 (golden run)
    versus one of the latest production runs, 5854. The single photoelectrons
    peaks line up well and the main amplitude is also similar. All the
    production runs from last night show the same behavior. I will keep on an
    eye on it to make sure the problem is gone.

    A copy of this log entry has been emailed to: bogdanw,jpchen,brads,jiang

    ADC Number 6